# The Expressive Capacity of State Space Models: A Formal Language Perspective

### v2

@@ B.4 Maintaining Counters
Theorem 23.
(Repeated from Theorem 5) Consider the languages Dyck-1, Shuffle-Dyck, n-ary Boolean Expressions, $a^{n}b^{n}$, $a^{n}b^{n}c^{n}$, and $a^{n}b^{n}c^{n}d^{n}$, as defined in Appendix C. Each of these is predictively modeled by an SSM.
@@
For Boolean Expressions: (here, $C$=1)
@@
For each of these mappings, we use Lemma 22 at $L=1$ to construct a one-layer SSMs that can, for each of the $C$ counters, distinguish the values 1,0,1$\leq$-1,0$\geq$ 1≤ - 1 , 0 , ≥ 1.
@@
Shuffle-$k$-Dyck is similar: EOS is allowed if and only if all counters are zero. An opening bracket is always allowed. A closing bracket is only allowed if the respective counter is $>0$.
@@ B.5 Bounded-Depth Dyck
The language $Dyck_{K,h}$ [28, 74] is given by the CFG with the nonterminals $\{S_{0},S_{1},\dots,S_{h-1},S_{h}\}$ and the following production rules:
@@
Theorem 25.
(Repeated from Theorem 6) The bounded-depth Dyck language $Dyck_{K,h}$ is predictively modeled by a two-layer SSM at finite precision, with $d=\mathcal{O}(h\log K)$.
@@ Appendix C Definitions of Languages
Here, we provide formal definitions of languages from the test suite based on Bhattamishra et al. [6]. Descriptions follow Bhattamishra et al. [6], and are included here for self-conrainedness. In all cases, our data generation setup is directly taken from [6].
@@ C.1 Regular Languages
Tomita Grammars. Used primarily as a benchmark language family for assessing sequence to sequence models [68], some of the languages in this family are star-free (with dot-depth of 1) and some non-star-free. All the regular languages of the family are defined on the alphabet $\Sigma=\{0,1\}$. Individual language definitions are available in Table 1.
$\bm{D_{n}}$. We follow the definition of [6] to define the $D_{n}$ family of star-free languages. In our experiments we only generate $D_{3}$, $D_{4}$, and $D_{1}2$ languages, given the proven equivalence of Tomita-2 and $D_{2}$. All the languages of the family are defined on the alphabet of $\Sigma=\{a,b\}$. $D_{n}=(aD_{n-1}b)^{*}$ has level $n$ in the dot-depth hierarchy.
@@ C.2 Counter Languages
We further use the family of Shuffle-k languages [51]. Shuffle-Dyck-k is defined in terms of $\Sigma=\{(_{1},)_{1},\dots,(_{k},)_{k}\}$. It is defined as the shuffle of $k$ Dyck-1 languages, each defined in terms of the alphabet $\Sigma_{i}=\{(_{i},)_{i}\}$ where $i=1,\dots,k$.
@@
This language is recognized by a counter automaton [19].
@@ 4 Yes strings without any 000’s substrings
5 No strings with an even number of 01’s and 10’s
@@ 7 Yes 0*1*0*1
Table 2: Accuracies on the counter Languages from the Bhattamishra et al. [6] test suite. Transformer results reported based on Bhattamishra et al. [6]. For Mamba, we report best settings (chosen based on inputs of length [1,50]) at 1 (Mamba1), 2 (Mamba2), 3 (Mamba3) layers. Results for the best-performing layer count, from the first two bins, are shown in Figure 3. On these languages, there is also a third bin.
@@
Table 3: Accuracies on the regular Languages from the Bhattamishra et al. [6] test suite - 1st half. Transformer results reported based on Bhattamishra et al. [6]. For Mamba, we report best settings (chosen based on inputs of length [1,50]) at 1 (Mamba1), 2 (Mamba2), 3 (Mamba3) layers. Results for the best-performing layer count are also shown in Figure 3.
@@
Table 4: Accuracies on the regular Languages from the Bhattamishra et al. [6] test suite - continued. Transformer results reported based on Bhattamishra et al. [6]. For Mamba, we report best settings (chosen based on inputs of length [1,50]) at 1 (Mamba1), 2 (Mamba2), 3 (Mamba3) layers. Results for the best-performing layer count are also shown in Figure 3.
@@ Appendix D Experimental Details
All experiments used the Mamba reference implementation [^7] https://github.com/state-spaces/mamba/blob/main/README.md. Unless stated otherwise, we followed the defaults given there ( $d_{state}=16$, $d_{conv}=4$, $expand=2$), as we found the default combination to work better than other options. We tuned $d_{model}$ for each language.
D.1 Bhattamishra et al. [6] test suite
@@
For all the languages, we use either the data prepared by Bhattamishra et al. [6] or—where not available—their data-generation scripts, allowing full comparability with results they reported for transformers. We used their official code and data release at https://github.com/satwik77/Transformer-Formal-Languages (last commit 48eea2e; MIT license). Training sets typically consist of 10K samples, with lengths varying between 1 to 50. There are two heldout bins: one with in-distribution lengths ([1,50]), and one testing length generalization (lengths [51,100]). The first one was used for hyperparameter optimization. Each bin typically contains around 2K samples. However for languages such as $a^{n}b^{n}$, where the number of positive examples in each bin was limited, all possible examples for that bin are included.
@@ D.2 FlipFlop
We obtained the dataset of Liu et al. [45] from their release, https://huggingface.co/datasets/synthseq/flipflop (MIT license).
Our setup corresponds to the deterministic (“clean”) mode in Liu et al. [45].
Matching Figure 2 in Liu et al. [45], we evaluated both with in-distribution data (matching the distribution of the training dataset) with $p_{i}=0.8,p_{w}=0.1,p_{r}=0.1$, and using an out of distribution sparse tail with $p_{i}=0.98,p_{w}=0.01,p_{r}=0.01$, where $p_{i},p_{w},p_{r}$ refer to the probabilities of that instruction appearing in input sequences.
@@
Following the evaluation criteria for LSTMs in Liu et al. [45], we compute the test every 100 training steps on our validation sets of choice, by randomly sampling around $10^{3}$ samples from each set in every evaluation cycle.
@@ D.3 Bounded Hierarchical Structure
We built on the official code and data release of Yao et al. [74] at https://github.com/princeton-nlp/dyck-transformer (last commit: 5d21fcf). We train a 2-layer Mamba and a 1-layer Mamba on $Dyck_{K,h}$ with $K=8$ and $h=10$. The training set and the validation set contains samples of lengths $\leq 700$, while the test set contains samples of lengths $700\leq n\leq 1400$. We train Mamba with a varying number of layers $l\in\{1,2\}$ and $d_{model}\in\{20,30,40,50,60,70,80,90,100\}$. We use the Adam optimizer with an initial learning rate of 0.01 or 0.001, using cross-entropy loss. After training for 100 epochs (with early stopping allowed in case of convergence), we select the learning rate with the better training performance.
@@ Appendix E Finite Precision Assumption
As described in Section 2, we adopt the finite precision notion used by Weiss et al. [70]: We allow an unbounded number of integer bits, but only $p$ fractional bits, where $p$ is a sufficiently large constant (e.g., $p=8$), independent of the length of the input.
@@
    Infinite precision Infinite precision allows any parameter and intermediate value to be an arbitrary number. Such a setting is unrealistic, as it would allow encoding arbitrary detail about the input into infinite precision [e.g. 63] and read these out with sufficiently powerful functions ($A$, $B$, $\phi$) in (5) – this would lead to the unrealistic conclusion that any function and language could be represented. For this reason, theoretical work has often adopted restricted precision notions.
@@
    Such a more restrictive definition would, however, mean that, for unbounded counting (Theorem 5), modeling is only possible up to a bound determined by the number of possible values—this is the one place where our results would be impacted. Indeed, we do observe that Mamba learns these counter languages on training lengths but struggles with length generalization. Transformers, on the other hand, can represent these languages with bounded activations (due to the constructions in [6]), and show strong length generalization.
@@
    An intermediary between infinite and finite precision is notions of precision where the number of allowed bits slowly increases with the input length, e.g., logarithmically. Such a setup has particularly been adopted for transformers [48], because a finite-precision assumption leads to very low expressivity in transformers. For SSMs, on the other hand, we find that finite precision assumptions are sufficient for showing a broad range of positive results.

### v3

@@ B.4 Maintaining Counters
Theorem 24 (Restated from Theorem 5).
The languages Dyck-1, Shuffle-Dyck, n-ary Boolean Expressions, $a^{n}b^{n}$, $a^{n}b^{n}c^{n}$, and $a^{n}b^{n}c^{n}d^{n}$, (defined in Appendix C) can each be predictively modeled by an SSM.
@@
For Boolean Expressions: (here, CC=1)
@@
For each of these mappings, we use Lemma 23 at $L=1$ to construct a one-layer SSMs that can, for each of the $C$ counters, distinguish the values $\leq-1,0,\geq 1$.
@@
Shuffle-$k$-Dyck is similar: EOS is allowed if and only if all counters are zero. An opening bracket is always allowed. A closing bracket is only allowed if the respective counter is >0>0.
@@ B.5 Bounded-Depth Dyck
The language $Dyck_{K,h}$ [hewitt2020rnns, yao-etal-2021-self] is given by the CFG with the nonterminals $\{S_{0},S_{1},\dots,S_{h-1},S_{h}\}$ and the following production rules:
@@
Theorem 26 (Restated from Theorem 6).
There is a two-layer SSM with $d=\mathcal{O}(h\log K)$ that predictively models $Dyck_{K,h}$ at all input lengths, at finite precision.
@@ Appendix C Definitions of Languages
Here, we provide formal definitions of languages from the test suite based on bhattamishra2020ability. Descriptions follow bhattamishra2020ability, and are included here for self-containment. In all cases, our data generation setup is directly taken from [bhattamishra2020ability].
@@ C.1 Regular Languages
Tomita Grammars. Used primarily as a benchmark language family for assessing sequence to sequence models [tomita1982dynamic], some of the languages in this family are star-free (with dot-depth of 1) and some non-star-free. All the regular languages of the family are defined on the alphabet $\Sigma=\{0,1\}$. Individual language definitions are available in Table 1.
$\bm{D_{n}}$. We follow the definition of bhattamishra2020ability to define the $D_{n}$ family of star-free languages. In our experiments, we only generate $D_{2}$, $D_{3}$, $D_{4}$, and $D_{12}$ languages; $D_{1}$ is equivalent to Tomita-2. All the languages of the family are defined on the alphabet of $\Sigma=\{a,b\}$. $D_{n}=(aD_{n-1}b)^{*}$ has level $n$ in the dot-depth hierarchy.
@@ C.2 Counter Languages
We further use the family of Shuffle-k languages [shuffle-dyck-def]. Shuffle-Dyck-k is defined in terms of $\Sigma=\{(_{1},)_{1},\dots,(_{k},)_{k}\}$. It is defined as the shuffle of $k$ Dyck-1 languages, each defined in terms of the alphabet $\Sigma_{i}=\{(_{i},)_{i}\}$ where $i=1,\dots,k$.
@@
This language is recognized by a counter automaton [Fischer1968].
@@ 4 Yes strings without any 000’s substrings
5 No strings of even length with an even number of 1’s
@@ 7 Yes 0*1*0*1
Table 2: Accuracies on the counter Languages from the bhattamishra2020ability test suite. Transformer results reported based on bhattamishra2020ability. For Mamba, we report best settings (chosen based on inputs of length [1,50]) at 1 (Mamba1), 2 (Mamba2), 3 (Mamba3) layers. Results for the best-performing layer count, from the first two bins, are shown in Figure 5. On these languages, there is also a third bin.
@@
Table 3: Accuracies on the regular Languages from the bhattamishra2020ability test suite - 1st half. Transformer results reported based on bhattamishra2020ability. For Mamba, we report best settings (chosen based on inputs of length [1,50]) at 1 (Mamba1), 2 (Mamba2), 3 (Mamba3) layers. Results for the best-performing layer count are also shown in Figure 5.
@@
Table 4: Accuracies on the regular Languages from the bhattamishra2020ability test suite - continued. Transformer results reported based on bhattamishra2020ability. For Mamba, we report best settings (chosen based on inputs of length [1,50]) at 1 (Mamba1), 2 (Mamba2), 3 (Mamba3) layers. Results for the best-performing layer count are also shown in Figure 5.
@@ Appendix D Experimental Details
All experiments used the Mamba reference implementation [^8] https://github.com/state-spaces/mamba/blob/main/README.md. xUnless stated otherwise, we followed the defaults given there ( $d_{state}=16$, $d_{conv}=4$, $expand=2$), as we found the default combination to work better than other options. We tuned $d_{model}$ for each language.
D.1 Test Suite from bhattamishra2020ability
@@
For all the languages, we use either the data prepared by bhattamishra2020ability or—where not available—their data-generation scripts, allowing full comparability with results they reported for transformers. We used their official code and data release at https://github.com/satwik77/Transformer-Formal-Languages (last commit 48eea2e; MIT license). Training sets typically consist of 10K samples, with lengths varying between 1 to 50. There are two heldout bins: one with in-distribution lengths ([1,50]), and one testing length generalization (lengths [51,100]). The first one was used for hyperparameter optimization. Each bin typically contains around 2K samples. However for languages such as $a^{n}b^{n}$, where the number of positive examples in each bin was limited, all possible examples for that bin are included.
@@ D.2 FlipFlop
We obtained the dataset of Liu2023FlipFlop from their release, https://huggingface.co/datasets/synthseq/flipflop (MIT license). Our setup corresponds to the deterministic (“clean”) mode in Liu2023FlipFlop. Matching Figure 2 in Liu2023FlipFlop, we evaluated both with in-distribution data (matching the distribution of the training dataset) with $p_{i}=0.8,p_{w}=0.1,p_{r}=0.1$, and using an out of distribution sparse tail with $p_{i}=0.98,p_{w}=0.01,p_{r}=0.01$, where $p_{i},p_{w},p_{r}$ refer to the probabilities of that instruction appearing in input sequences.
@@
Following the evaluation criteria for LSTMs in Liu2023FlipFlop, we compute the test every 100 training steps on our validation sets of choice, by randomly sampling around $10^{3}$ samples from each set in every evaluation cycle.
@@ D.3 Bounded Hierarchical Structure
We built on the official code and data release of yao-etal-2021-self at https://github.com/princeton-nlp/dyck-transformer (last commit: 5d21fcf). We train a 2-layer Mamba and a 1-layer Mamba on $Dyck_{K,h}$ with $K=8$ and $h=10$. The training set and the validation set contains samples of lengths $\leq 700$, while the test set contains samples of lengths $700\leq n\leq 1400$. We train Mamba with a varying number of layers $l\in\{1,2\}$ and $d_{model}\in\{20,30,40,50,60,70,80,90,100\}$. We use the Adam optimizer with an initial learning rate of 0.01 or 0.001, using cross-entropy loss. After training for 100 epochs (with early stopping allowed in case of convergence), we select the learning rate with the better training performance.
@@ Appendix E Finite Precision Assumption
As described in Section 2.1, we adopt the finite precision notion used by weiss2018practical: We allow an unbounded number of integer bits, but only $p$ fractional bits, where $p$ is a sufficiently large constant (e.g., $p=8$), independent of the length of the input.
@@
    Infinite precision Infinite precision allows any parameter and intermediate value to be an arbitrary number. Such a setting is unrealistic, as it would allow encoding arbitrary detail about the input into infinite precision [e.g. siegelmann2012neural] and read these out with sufficiently powerful functions ($A$, $B$, $\phi$) in (4) – this would lead to the unrealistic conclusion that any function and language could be represented. For this reason, theoretical work has often adopted restricted precision notions.
@@
    Such a more restrictive definition would, however, mean that, for unbounded counting (Theorem 5), modeling is only possible up to a bound determined by the number of possible values—this is the one place where our results would be impacted. Indeed, we do observe that Mamba learns these counter languages on training lengths but struggles with length generalization. Transformers, on the other hand, can represent these languages with bounded activations (due to the constructions in bhattamishra2020ability), and show strong length generalization.
@@
    An intermediary between infinite and finite precision is notions of precision where the number of allowed bits slowly increases with the input length, e.g., logarithmically. Such a setup has particularly been adopted for transformers [merrill2023logic], because a finite-precision assumption leads to very low expressivity in transformers. For SSMs, on the other hand, we find that finite precision assumptions are sufficient for showing a broad range of positive results.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction summarise the theoretical and empirical results. We took to only include well-supported claims.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The Discussion section includes a paragraph on Limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Every theorem is proven formally in the appendix. Assumptions about the SSM architecture are stated formally.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Besides describing experimental details, we have publicly released our codebase on GitHub and added a link to it in the main paper with instructions on how to run our experiments, ensuring that our findings can be reproduced.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Yes, we have publicly released our codebase on GitHub and added a link to it in the main paper with instructions on how to run our experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have specified the training and test details in a dedicated appendix section. Our experiments use freely available dataset or dataset generating scripts from previous studies, which we cite. Our uploaded code base ensures that all settings are reproducible.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: In the empirical results relevant to supporting our positive theoretical results, the accuracy attained by Mamba is either very close to 100% or far from 100%. As the test sets include hundreds or thousands of data points, error bars would not add substantial further information.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided information in the appendix section on experimental setup.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have verified compliance with the NeurIPS Code of Ethics,
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The paper studies foundational properties of neural network architectures, and we foresee no positive or negative societal impact of the results.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We are not releasing data or models, and foresee no risk of misues of our theoretical results.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The paper cites the original papers, and provides details in the Appendix section on experimental details.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper releases no new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper involves no crowdsourcing nor human subjects research.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper involves no crowdsourcing nor human subjects research.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.